### v1

# External charged debris in a flowing plasma : charge fluctuation induced complexity

Abstract
In this work, we investigate the response of a flowing $e$-$i$ plasma to embedded external charged debris, focusing on the periodic debris charge fluctuations that can trigger complex phenomena such as chaos and nonlinear Landau damping. We employ both kinetic and fluid simulations to analyse the plasma response to the time-dependent debris charge. Our findings indicate that the nature of nonlinear response can be considerably different for fluctuating positively charged external debris from a negatively charged debris. The simulations show that the debris charge fluctuation causes damping of the ion-acoustic wave as the debris velocity nears the ion-acoustic speed through nonlinear Landau damping and wave-wave interactions. We also present a theoretical framework to support the simulation findings. Our findings provide critical insights into debris-plasma interactions which may be useful in applications involving space debris management.
keywords:
Plasma-Debris Interaction , Charge Fluctuation , Chaos , Nonlinear Landau damping
journal: Chaos, Solitons & Fractals
[…]
organization=Affiliation: Physics Department, Gauhati University,addressline=Gauhati University, city=Guwahati, postcode=781014, state=Assam, country=India
1 Introduction
Impurities or external entities in a typical electron-ion plasma is not a new subject and the arena of scientific research involving such “foreign” entities in plasma environments is dotted with numerous studies, which essentially began with Irving Langmuir’s pioneering work on plasma sheath in the 1920s and 1930s [1, 2, 3, 4, 5, 6]. One of the most distinctive branches which stemmed out of such plasma physics research is the so-called “dusty plasma”, which has become one of the core research areas in plasma physics [7, 8, 9, 10]. Another such area, which came to the notice of the scientific community is the interaction of the plasma with charged external or foreign bodies, when the importance of scientific studies of such interactions was realised during the early days of space exploration, in the 1960s, when several space missions began studying how satellite and spacecraft interact with space plasma environments [11, 12, 13, 14]. In recent years, research in this area has really accelerated, especially after the subject of space debris in relation to the safety of the satellites and space stations in low Earth orbits (LEOs) has been recognised as an important research area. Usually, space debris at LEO are detected with satellite and ground-based sensors operating at optical and radio ranges and these methods are not effective in detecting smaller debris (size $\lesssim 1\,{\rm cm}$) [15], though NASA’s Haystack Ultrawideband Satellite Imaging Radar (HUSIR) has been able to provide a map of LEO debris down to $\sim 5.5\,{\rm mm}$ [16]. Recent techniques, based on detection of plasma oscillations excited by these smaller debris, have been tested successfully to detect some of these debris [17]. In essence, the number of proven and effective methods to detect these smaller debris, is by no means complete and we continue to look toward understanding different physical manifestations of plasma-debris interaction. And this work is an effort to understand the intricacies of these interactions by allowing more realistic physical conditions.
In recent times, our understanding about plasma-debris interaction has been steadily progressing due to several theoretical, experimental, and numerical studies [18, 19, 20, 21, 22, 23]. It is now definitively proved both theoretically and through numerical simulations that the depending on the nature of the charged debris, different nonlinear waves in the ion-acoustic regime can get excited by moving debris [21, 23]. While we usually see dispersive shock waves due to a positively charged external debris, pinned solitons are formed due to phase space trapping through ion-ion counter streaming instability (IICSI) due to a negatively charged debris [23]. Some very recent laboratory and numerical simulation studies have looked into the plasma-debris interaction in multi-ion plasma [22]. Several other studies have focused on shape, dimension, and speed of the debris on the propagation of these solitons theoretically and with experiments [24, 20], as well as bending and acceleration [25], damping (both collisional and Landau) [26] of the excited structures.
In this work, we plan to explore a yet unexplored area related to plasma-debris interaction by allowing the charge in the external debris to vary over time along with the presence of a stochastic component. There are good reasons as to why the charge on external debris should vary. Some seemingly plausible causes are periodic photoemission due to eclipse transitions (especially debris in the LEO), passage of the debris through the regions of density and/or temperature gradient of the ambient plasma, plasma wave interactions, intrinsic forcing due to spin or tumbling motion of the debris, and external modulations viz. during spacecraft remediation processes, among others. Out of these scenarios, the most important one is perhaps the debris charge modulation due to plasma wave interactions as this is an inherent process and possibly cannot be avoided. Debris moving at or near ion-acoustic speed is expected to make the charge on debris oscillate in time at ion-plasma frequency $(\omega_{pi})$, through ion-acoustic wake oscillations [27]. As an ion-acoustic wave (IAW) passes over the debris, it periodically alters the ion and electron densities. Since the ion and electron currents depend sensitively on these densities, the wave creates an imbalance in the ion and electron fluxes to the debris surface. This imbalance, in turn, disturbs the equilibrium floating potential of the debris and the debris charge adjusts dynamically to restore the balance, leading to a periodically oscillating debris charge. So, basically IAWs excited by the external charged debris, itself can cause the periodic fluctuation of the debris charge at frequencies $\sim\omega_{pi}$, which is very much inherent to the debris-plasma interaction scenario. This process is very much similar to the process of dust-charge fluctuation in a dusty plasma [27, 28]. This inherent charge fluctuation of the debris can be thought to be a forcing term, which can result in parametric coupling and can give rise to a wide range of nonlinear phenomena including chaos, sheath instability, nonlinear Landau damping [29] (NLLD) and can also affect the local turbulence. In this analysis, we primarily focus on two effects – chaos and NLLD, the signatures of which are most prominent. While detecting chaos is pretty straight forward, the same is not true for NLLD, which requires more than one confirmation methods. We believe that the periodic oscillation near $\omega_{pi}$ can trigger NLLD, which can then limit the wave amplitude, providing a feedback mechanism on the charging process and may eventually affect the nonlinear structures [30]. On the other hand, as chaotic oscillations can imply the breakdown of coherent wave-particle energy exchange, it can act as a probe for the diagnostic tool the wave-particle interaction limit [31, 32].
The paper is organized as follows. In Section II, we present the results of kinetic and fluid simulations with the help of our h-PIC-MMC and mFCT codes, respectively. In Section III, we provide a theoretical analysis of the results obtained in Section II. We provide an additional analysis of the results obtained in earlier sections, in terms of power spectrum analysis in Section IV. In Section V, we conclude. We also provide a brief analysis of the effect of fluctuation of debris charge on turbulent dynamics.
2 Simulation results
2.1 The premise
Though technically it should be possible to model a self-consistent periodic fluctuation of charge in an external debris embedded in a plasma, in reality it is much harder to model it mathematically as it would naturally require calculation of collision cross-sections to exactly determine the electron and ion currents to the debris, which may not obey the much used orbit motion limited (OML) theory, usually used to calculate the currents. It may be further complicated by extended and arbitrary shape of the debris. Besides, a self-consistent fluctuation will necessarily require a coupling a circuit-like differential equations with the partial differential equations of the fluid model for the plasma. So, for a fluid simulation, we shall require a charging equation like
$\frac{dQ_{{\rm d}}}{dt}=I_{i}(t)+I_{e}(t)+I_{{\rm wave}}(t)+\cdots,$ (1)
to be coupled to the plasma fluid equations, where $Q_{d}$ is the charge on the debris and $I_{i,e,{\rm wave}}$ are the currents to the debris due to plasma ions, electrons, and the IAW. For a fluid simulation, the most challenging part would be to model the “wave” current. On the other hand, a PIC simulation may naturally model the $I_{{\rm wave}}$ term through kinetic effects, but it does not natively compute the above charging equation dynamically and would require some kind of approximation with a sheath and potential model around the debris. Besides, it would also require much higher grid resolution for realistic calculation of these currents to the debris. As a result, one would ideally need a hybrid approach of fluid and PIC methods to model these fluctuations self-consistently.
In order to avoid these complexities, we have chosen to model the periodic fluctuation of charge on external debris through an “external forcing” equation for the debris charge density $\rho_{{\rm ext}}$
$\rho_{{\rm ext}}(t)=\rho_{0}{\cal F}(t),$ (2)
where ${\cal F}(t)$ is the time-dependent part. This model can be seamlessly incorporated into a fluid as well as a PIC model and the results can be very efficiently compared. Besides, the controllable external forcing-like equation allows us to have a finer control on the fluctuation parameters and helps us isolate different nonlinear phenomena through controlled tuning. In what follows, we shall present results from two different simulation methods – PIC simulation and fluid simulation with a flux-corrected transport solver, which justifies these observations.
2.2 The plasma model
The 1-D $e$-$i$ plasma system with Boltzmannian electrons is modeled with the following equations with an exclusive time-dependent external charge density for the debris,
$\frac{\partial n_{i}}{\partial t}+\frac{\partial}{\partial x}(n_{i}v_{i})$ $=$ $0,$ (3)
$\frac{\partial v_{i}}{\partial t}+v_{i}\frac{\partial v_{i}}{\partial x}$ $=$ $-\frac{1}{m_{i}n_{i}}\frac{\partial p_{i}}{\partial x}-\frac{e}{m_{i}}\frac{\partial\phi}{\partial x},$ (4)
$n_{e}$ $=$ $n_{0}\exp\left(\frac{e\phi}{T_{e}}\right),$ (5)
$\epsilon_{0}\frac{\partial^{2}\phi}{\partial x^{2}}$ $=$ $e(n_{e}-n_{i})-\rho_{{\rm ext}}(t;x-v_{d}t),$ (6)
where the symbols have their usual meanings. The pressure for ions is treated as polytropic
$p_{i}\propto n_{i}^{\gamma},$ (7)
where $\gamma$ is the ratio of specific heats. The last term in the Poisson equation is the term due to external charged debris which is moving with respect to the bulk plasma at a constant velocity $v_{d}$. Note that depending on whether the debris charged positively or negatively, $\rho_{{\rm ext}}\lessgtr 0$. The normalized model can be written as
$\frac{\partial n_{i}}{\partial t}+\frac{\partial}{\partial x}(n_{i}v_{i})$ $=$ $0,$ (8)
$\frac{\partial v_{i}}{\partial t}+v_{i}\frac{\partial v_{i}}{\partial x}+\gamma\sigma n_{i}^{\gamma-2}\frac{\partial n_{i}}{\partial x}$ $=$ $-\frac{\partial\phi}{\partial x},$ (9)
$n_{e}$ $=$ $e^{\phi},$ (10)
$\frac{\partial^{2}\phi}{\partial x^{2}}$ $=$ $n_{e}-n_{i}-\rho_{{\rm ext}}(t;x-v_{d}t),$ (11)
where the densities are normalised by the equilibrium plasma density $n_{0}$, length is normalized by electron Debye length $\lambda_{D}$, time is normalized by the inverse of ion-plasma frequency $\omega_{pi}$, plasma potential $\phi$ is normalized by $(T_{e}/e)$. The temperature is expressed in energy units and and velocity is normalized by the ion-acoustic speed $c_{s}=\sqrt{T_{e}/m_{i}}$. The quantity $\sigma$ is the ratio of ion to electron temperature and is usually $\ll 1$. In the above equations, the external charge density $\rho_{{\rm ext}}$ is also normalized accordingly.
We consider the external charge density to be either positive definite or negative definite, so that the charge in particular debris remains same (either positive or negative) while oscillating in time.
2.3 Kinetic model - PIC simulation
In this section, we shall carry out a PIC simulation of the above-mentioned $e$-$i$ plasma with a time-varying external charged debris. The simulation is carried out with our well-tested hybird-PIC-MCC code [33, 34, 35, 23], which has already been successfully applied to several electrostatic problems including plasmas with secondary electron emission, kinetic dust-ion-acoustic wave with ion-Landau damping effect, and $e$-$i$ dusty plasma with external charged debris.
The plasma is assumed to have an $e$-$i$ number density $n_{e,i}\sim 10^{16}\,{\rm m}^{-3}$, with electron temperature $T_{e}\sim 1\,{\rm eV}$ and ion temperature $T_{i}\sim 0.01\,{\rm eV}$. The simulation box length $L\sim 140\lambda_{D}\sim 0.01\,{\rm m}$. The code can handle any arbitrary-shaped external charge distribution of arbitrary strength which can be either positive or negative with time-varying charge density. We can either make the debris move across the plasma or vice versa – the results being same in either case. Computationally, as mentioned before, we model our time-varying charge density with the following function
$\rho_{{\rm ext}}(t)=\rho_{0}[1+\tanh\{\Delta\sin(\nu t)\}]+\rho_{{\rm noise}}(t),$ (12)
which is a symmetric (Gaussian-like) charge distribution with $\Delta$ as the flatness parameter of the time variation and $\nu$ as the oscillation frequency of the charge distribution. The larger the value of $\Delta$ is, the steeper is the variation. Note that the $\rho_{{\rm noise}}(t)$ part denotes a stochastic part of $\rho_{{\rm ext}}(t)$, which adds a degree of random-ness to the oscillation. The variable data $(v_{i},n_{i})$ are extracted across the simulation domain as independent time series, which are then examined with Wolf’s algorithm for possible chaotic nature. It is now well known that Wolf’s algorithm [36] applies Taken’s theorem [37] to construct an $m$-dimensional phase space from a scalar time series $\{x_{t}\}$,
$\bm{X}_{t}=\left[x_{t},x_{t+\tau},x_{t+2\tau},\cdots,x_{t+(m_{{\rm PIC}}-1)\tau}\right],$ (13)
with $m_{{\rm PIC}}$ as the embedding dimension and $\tau$ as the time delay. The algorithm continuously measures the divergence between two nearest neighbor (forward in time)
$\delta(t)=\left\|\bm{X}_{t}-\bm{X}^{\prime}_{t}\right\|,$ (14)
where $\bm{X}^{\prime}_{t}$ is $\bm{X}_{t}$’s nearest neighbor (in time). The so-called Lyapunov exponent [38], which is nothing but the exponential divergence of the nearest trajectories, is then calculated over valid divergence segments as
$\lambda_{j}=\frac{1}{\Delta t}\,\ln\left[\frac{\delta(t+\Delta t)}{\delta(t)}\right],$ (15)
where $\Delta t$ time difference taken to compute the divergence. The maximal Lyapunov exponent is then calculated by
$\lambda=\frac{1}{N}\sum_{j=1}^{N}\lambda_{j},$ (16)
where $N$ is the number of valid divergence segments. Any positive $\lambda$ indicates chaos and the further $\lambda$ is away from 00, the stronger is the chaos. A typical rule of thumb is to have $\lambda\gtrsim 0.01$ for a definitive proof of chaos. In our case, we have used an embedding dimension $m_{{\rm PIC}}=4$ for our chaos test.
Figure 1: Maximal Lyapunov exponent $\lambda$, shown in the $x$-$\nu$ plane for a positively charged external debris (shown as the orange strip). The blue and red points indicate values of $\lambda$ with and without stochastic fluctuation. The results are at time $\sim 20\tau_{pi}$.
Figure 2: Maximal Lyapunov exponent $\lambda$, shown in the $x$-$\nu$ plane for a negatively charged external debris (shown as the orange strip). The blue and red points indicate values of $\lambda$ with and without stochastic fluctuation as before. The results are at time $\sim 20\tau_{pi}$.
The result of the simulation is shown in Figs.1 and 2, where we have shown the scatter plots of $\lambda$ in a $x$-$\nu$ plane. The position of the debris is indicated by an orange-colored strip in the panels. The position is expressed in terms of electron Debye length $\lambda_{D}$ and the frequency $\nu$ is expressed in terms of ion plasma frequency $\omega_{pi}$. In order to help the perspective visualization of the value of $\lambda$, we have added “vertical drop-lines” from every point to the $z=0$ baseline so that the length of the ‘drop-line’ indicates the value of $\lambda$. All simulation results are at time $\sim 20\tau_{pi}$, where $\tau_{pi}$ is the plasma period $=\omega_{pi}^{-1}$. The results for positively charged debris is shown in Fig.1 and the negatively charged results are shown in Fig.2. As we can see from Fig.1 that for positively charged debris, increase in debris velocity definitely induces quite stronger chaos in the oscillation, specifically at the site of the debris. The average strength of the $\lambda$ in panel (a) is almost in the boundary line of chaos $\sim 0.01$, which is without any relative motion between the plasma and external debris $(v_{d}=0)$. This should be contrasted with those in panels (b), (c), and (d), where we can see a definitive built up of chaotic oscillations. It should also be noted that the site of chaotic oscillation is mostly confined to the debris site. Besides, the stochasticity of fluctuation is not seen to have any systematic and correlated effect on $\lambda$, a situation which we shall explain later in Section 3.2.4.
In contrast to the positively charged debris, we can see a completely different scenario in case of negatively charged debris from Fig.2. In this case, we can see that the strength of the chaos is quite weak as compared to the chaos induced by fluctuating positively charged debris. Besides, the chaotic points are almost evenly distributed across the $x$-$\nu$ plane as opposed to the earlier case where the chaos is mostly confined to the debris site. As far as the effect of stochastic fluctuation is concerned, it is similar to that of positively charged debris.
The corresponding ion phase space and ion density plots are shown in Fig.3 for positive debris and in Fig.4 for negative debris. All frames of Figs.3 and 4 are drawn in the rest frame of the debris, which lies in the middle of the simulation window denoted by a dashed vertical line in each panel. The ion velocity is normalized to the ion-acoustic speed and the ion density is normalized by its equilibrium value. The length is normalized by electron Debye length $\lambda_{D}$. All these plots are without any stochastic fluctuation component. What we can see from the plots is that, the periodic charge fluctuation of the external debris has a minimal effect when the debris charge is negative, also manifested by weak chaos, as shown before.
Figure 3: Ion phase space plots and ion density plots are shown in panels (a), (b) and density (c), (d) for positively charged external debris. The left hand plots (black colored) are without any fluctuation of the debris charge and the right hand plots (red colored) are for periodic fluctuation of the debris charge. The value of the fluctuation frequency $\nu$ is given in each panel.
Figure 4: Ion phase space plots and ion density plots for negatively charged external debris. All other particulars are same as in Fig.3. The value of the fluctuation frequency $\nu$ is given in each panel.
2.4 Fluid model – FCT simulation
In this section, we shall try to model the fluctuation-induced chaos from a fluid perspective with flux-corrected transport simulation [21, 22]. Here, we use a multi-fluid FCT (mFCT) code based on Boris’s original algorithm [39] with Zalesak’s flux limiter [40], which we use to solve Eqs.(3-6) with time-varying $\rho_{{\rm ext}}$. This code has also been extensively used to study debris-induced plasma wave phenomena [22].
The FCT formalism requires the hydrodynamic equations Eqs.(3-6) to be put in the form of a generalized continuity equation,
$\frac{\partial f}{\partial t}=-\frac{\partial}{\partial x}(fv)+\frac{\partial s}{\partial x},$ (17)
where $f=(n_{i},v_{i})$ is the physical quantity to be solved, $(fv)$ is the corresponding flux, and $s$ is the source term. The external charged debris is entered through the Poisson equation Eq.(6) as before which is to be solved each time step along with Eq.(17). The external charged debris is modeled with the same profile as used in the PIC simulation, given by Eq.(12).
We now present a full-spectrum Lyapunov exponent (LE) $\lambda_{k}$ analysis of the velocity profile on the debris site, as obtained from FCT simulation. The algorithm to calculate Lyapunov exponent is based on Wolf’s algorithm [36] for an embedding dimension $m_{{\rm FCT}}=6$.
$\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{m},\quad m_{{\rm FCT}}=6,$ (18)
for a scalar time series $v_{i}$, at the site of the debris, resulted out of the fluid simulation. We start by constructing the delay vectors
$\vec{f}_{j}=\left[f(t_{j}),f(t_{j}+\delta t),\dots,f(t_{j}+(m_{{\rm FCT}}-1)\delta t)\right]\,{\rm in}\,\mathbb{R}^{6},$ (19)
required for phase-space reconstruction. The embedding dimension $m_{{\rm FCT}}=6$ is further verified through Kaplan-Yorke (KY) dimensional analysis [41] of the embedding dimension,
$D_{{\rm KY}}=j+\frac{\sum_{k=1}^{j}\lambda_{k}}{|\lambda_{k+1}|},\quad\sum_{k=1}^{j}\lambda_{k}>0,$ (20)
where the summation is over all non-zero $\lambda_{k}$. This yields KY fractal dimension $D_{{\rm KY}}\approx 5$ for our $\lambda_{k}$s which conforms to the general guideline $m_{{\rm FCT}}>D_{{\rm KY}}$, an over-conservation as compared to Taken’s theorem [37] that in general requires $m_{{\rm FCT}}>2D_{{\rm KY}}$. While calculating the Lyapunov spectrum, we have used a Theiler window [38] to avoid temporal correlation and tracked the divergence of the nearby trajectories using a linearized approximation. A Gram-Schmidt orthonormalization (i.e. QR decomposition) [42] is applied in each time step to the evolving tangent vectors and the LEs are calculated finally by
$\lambda_{k}=\frac{1}{T}\sum_{k=1}^{N}\log R_{kk}^{(p)}.$ (21)
In the above expression, $R$ is the upper-triangular matrix as obtained from the QR decomposition, $p$ is the index of time step at which the decomposition is carried out, $N$ is total number of such steps, and $T$ is the physical time $T=\sum_{k=1}^{N}\Delta t_{k}$ with $\Delta t$ being time interval (of the time series) between each decomposition step.
Note that the reason for the embedding dimension for FCT simulation $(m_{{\rm FCT}}=6)$ being higher than that for PIC simulation $(m_{{\rm PIC}}=4)$ can be understood by considering the presence of “noise” in PIC simulation, which is almost absent in FCT simulation as it numerically solves the model equations exactly. On the other hand, the “noise” in PIC simulation scales as inverse of the square root of total number of computational particles [43]
$\textrm{PIC Noise}\propto\frac{1}{\sqrt{N_{p}}},$ (22)
where $N_{p}$ is the total number of computational particles used in the PIC simulation. As computational cost per PIC cycle is $\propto N_{p}$, a decrease in “noise” by a factor of $2$ increases computational cost by a factor of $4$, making the option of increasing $N_{p}$ not a viable one. Naturally, the FCT simulation can detect more fine-scale fluctuations whereas PIC simulation picks up more coarse-grained fluctuations which results in a low-dimensional dynamical system for PIC simulation.
The results of this analysis is shown in Fig.5, which shows the full spectrum LE. In the figure, the axes are the three dominant LEs $(\lambda_{1,2,3})$ and the red and blue circles respectively denotes the cases $\rho_{{\rm ext}}>0$ and $<0$. Apart from the fact that all $\lambda_{1,2,3}>0$ which signifies chaos, the accumulation of the red circles at the upper portion of the figure indicate a considerably stronger chaos for $\rho_{{\rm ext}}>0$ than $<0$. The four cases corresponding to four circles in the figure (red and blue) are for different debris velocities $v_{d}=0.2,0.5,1.0,$ and $1.5$ (The numerical code used to calculate the full-spectrum LE is available).
Figure 5: Full spectrum LEs calculated from the time series obtained from the FCT simulation.
Figure 6: Suppression of density oscillations with periodic charge fluctuation of the negatively charged debris. The top panels (a), (b) are obtained from PIC simulation while the bottom panels (c), (d) are obtained from FCT simulation. The panels on the left are without any charge fluctuation and those on the right are with charge fluctuation.
2.5 Nonlinear Landau damping?
We now present a combined result of PIC and FCT simulations which points out a phenomenon similar to what is known as “nonlinear Landau damping” [29]. In both cases, we observe a suppression of oscillations when $v_{d}\sim c_{i}$, where $c_{i}=\sqrt{T_{e}/m_{i}}$ is the phase velocity of the ion-acoustic oscillations. However, the fundamental physics behind this observation is slightly different in both cases.
Figure 7: Flattening of the ion VDF due to fluctuation. While the top panels (a) and (b) display respectively the VDFs calculated from the PIC simulation without $(\nu=0)$ and with $(\nu\sim 1)$ fluctuation, the bottom panels (c) and (d) show the numerically calculated FWHMs for the corresponding VDFs. The panel (d) on the right clearly shows that the VDF is flattening due to fluctuation.
Consider an external charged debris which is moving with a velocity $v_{d}$ with a potential profile $\phi(x-v_{d}t)$, modulated by a time-varying charge $Q(t)$. The plasma responds to this debris field through ion-acoustic oscillations with a phase velocity $c_{i}$. So, when $v_{d}\sim c_{i}$, we have a resonant energy transfer between the background plasma and the charged debris. However, the oscillating charge of the debris is going to introduce another frequency scale $\sim\nu$. As a result, the Doppler-shifted frequency of the oscillation (as seen by the debris) is given by
$\omega_{{\rm forcing}}=\nu+kv_{d}.$ (23)
Thus, we have a resonance condition, when $\omega_{{\rm forcing}}\simeq\omega_{{\rm IA}}=kc_{i}$, or
$v_{d}\simeq c_{i}-\frac{\nu}{k}.$ (24)
We can expect a strong wake when a resonance occurs and destructive interference when it becomes out of phase,
$|\nu+k(v_{d}-c_{i})|\gtrsim 0,$ (25)
reaching a maximum of $\pi$ when they become completely out of phase by a turnover time of $\tau_{\pi}$
$\tau_{\pi}|\nu+k(v_{d}-c_{i})|\sim\pi.$ (26)
It is then more likely that ‘anti-resonance’ occurs via relations (25) and (26) rather than a resonance condition occurs, which has to be exactly via relation (24). So, we expect a suppression of oscillations to occur, on average, due to the periodic variation of debris charge. This is the reason behind the observation of suppression of oscillation seen in the FCT simulation results. However, as the PIC simulation is expected to also incorporate the effect of NLLD, we perform an additional check on the ion velocity distribution function (VDF) as obtained from PIC simulation. Though there are no definitive proofs of NLLD in general, for collision-less electrostatic plasma such as ours, a flattening of the velocity distribution can be attributed to NLLD.
This effect is shown in Fig.6 for negatively charged debris, both as obtained from PIC and FCT simulations. However, this resonant suppression of oscillation occurs at $v_{d}\lesssim c_{i}$ for PIC simulation and at $v_{d}\gtrsim c_{i}$ for FCT simulation as can be seen from the figure. This discrepancy can be explained from the fact that the PIC simulation captures the full kinetic response including wave-particle resonance and phase space trapping while the FCT simulation captures only the macroscopic fluid-like wave-wave response and lacks phase mixing. In Fig.7, we have shown the ion VDF obtained from PIC simulation data with and without debris charge fluctuation, which clearly show the flattening of the VDF with fluctuation. The suppression of oscillations combined with the definitive proof of flattening of the ion VDF in PIC simulation strongly supports the fact that oscillations undergo NLLD due to charge fluctuation in the case of negatively charged debris. We note here that this flattening of ion VDF is not seen in case of fluctuating positively charged debris.
3 Theoretical analysis
3.1 Torus breakdown of a chaotic plasma model
In this section, we try to understand the theory behind the chaotic oscillation observed in the simulations and try to understand some of the observations related to stochastic fluctuations and why the properties of chaos for positively and negatively charged debris are fundamentally different. In particular, we shall discuss how the fluctuation of charge density of a debris may force the system to undergo torus breakdown [44, 31, 45] and become chaotic.
In order to facilitate a dynamical analysis, we make a coordinate transformation to the moving frame of the debris with the introduction of a scaled time $\tau=t-x/v_{d}$, so that in terms of the scaled variable we have $\partial/\partial t\to\partial/\partial\tau$ and $\partial/\partial x\to-v_{d}^{-1}\partial/\partial\tau$ [35]. With this transformation, our plasma model represented by Eqs.(3-6) becomes
$\frac{\partial n_{i}}{\partial\tau}-\frac{1}{v_{d}}\frac{\partial}{\partial\tau}(n_{i}v_{i})$ $=$ $0,$ (27)
$\frac{\partial v_{i}}{\partial\tau}-\frac{v_{i}}{v_{d}}\frac{\partial v_{i}}{\partial\tau}-\frac{\gamma\sigma}{v_{d}}n_{i}^{\gamma-2}\frac{\partial n_{i}}{\partial\tau}$ $=$ $\frac{1}{v_{d}}\frac{\partial\phi}{\partial\tau},$ (28)
$\frac{1}{v_{d}^{2}}\frac{\partial^{2}\phi}{\partial\tau^{2}}$ $=$ $n_{e}-n_{i}-\rho_{{\rm ext}}(\tau).$ (29)
Integration of Eq.(27) gives us
$n_{i}=\frac{v_{d}-v_{0}}{v_{d}-v_{i}},$ (30)
where $v_{0}$ is the velocity of the ions at infinity (or far away from the debris). In the above integration, we have also used the boundary condition that at $\infty$, $n_{i}\to 1$. Note that for the above formalism to work, we must have $v_{d}\lessgtr(v_{0},v_{i})$ or else the density would be negative, which is unphysical. Without loss of any generality we can assume $\gamma=2$ which helps us to integrate Eq.(28) to have
$\phi=\frac{1}{2}\left(v_{0}^{2}-v_{i}^{2}\right)-v_{d}(v_{0}-v_{i})+2\sigma\left(1-\frac{v_{d}-v_{0}}{v_{d}-v_{i}}\right),$ (31)
where we have used the expression given by Eq.(30). Inserting the above expressions in Poisson equation, we have the following nonlinear second order differential equation in $v_{i}$,
$\ddot{v}_{i}A+\dot{v}_{i}^{2}B+C=0,$ (32)
where the ‘$\,\dot{}\,$’ refers to derivative in $\tau$ and
$A$ $=$ $\frac{1}{v_{d}^{2}}\left(\mu_{i}-2\sigma\frac{\mu_{0}}{\mu_{i}^{2}}\right),$ (33)
$B$ $=$ $-\frac{1}{v_{d}^{2}}\left(1+4\sigma\frac{\mu_{0}}{\mu_{i}^{3}}\right),$ (34)
$C$ $=$ $\rho_{{\rm ext}}(\tau)+\frac{\mu_{0}}{\mu_{i}}-e^{\phi}$ (35)
where $\mu_{i}=v_{d}-v_{i}$ and $\mu_{0}=v_{d}-v_{0}$. Note that Eq.(32) is a highly nonlinear non-autonomous equation.
Figure 8: Poincaré plots of one transition phase from periodic to a chaotic state through torus breakdown. The initial (periodic) state is shown at the left top and the final chaotic state is shown at the right top panels. The arrow and the blown up transition phases are shown in the lower smaller panels.
Figure 9: Power spectrum density at transition and maximal Lyapunov exponent at the final chaotic state.
3.1.1 Torus breakdown
As we have mentioned that the system is highly nonlinear, it can undergo a chaotic transition with the variation of different control parameters. The fluctuation frequency $\nu$ of the charge density of the external debris and the debris velocity $v_{d}$ are two such parameters. As we shall observe, the system makes several quick transitions from quasi-periodic to chaotic states and vice versa when the control parameters are varied. So, there are several non-chaotic states of the system interspaced with numerous stages when the system undergoes torus breakdown and becomes chaotic. These breakdowns can be very quick and dramatic and do not follow a fixed pattern – a hallmark of torus breakdown [44, 31, 45]. One such transition phase is shown in Fig.8 through the Poincaré plots.
The charge fluctuation of the external debris is modeled by periodically switching the debris term on and off with a frequency $\nu$. Toward this, the external debris term is expressed in a periodic top-hat function as mentioned before
$\rho_{{\rm ext}}(\tau)=\rho_{0}[1+\tanh\{\Delta\sin(\nu\tau)\}],$ (36)
where $\rho_{0}$ denotes of the strength of the charge density and $\Delta$ controls the steepness of the top-hat function which is typically $\sim 8-10$. As our system is normalized, $\rho_{0}$ is typically set to $\sim 0.5$, which is about $50\%$ of the background plasma density. The fluctuation frequency $\nu$ is set to $\sim 1-5$ or about five ion-plasma periods. In Fig.8, we have shown torus breakdown of the system where, in the top row, the Poincaré map of the initial deformed torus (quasi-periodic) is shown, which makes a final transition to a chaotic state on the right when the breakdown is complete. The successive transition from the initial state to the final state is depicted in the smaller panels at the lower rows for shaded region marked in the bigger panels. The whole transition occurs within the interval of the debris velocity $v_{d}\in[1.596,1.6]$. The torus breakdown is further confirmed by the power spectrum density $P$ of the oscillation, shown in the left panel of Fig.9, where one can see the numerous random frequencies dominated by some prominent frequencies. The final chaotic state is also confirmed by the maximal Lyapunov exponent $\lambda$ shown in the right panel of Fig.9, which settles down at about $\sim 0.03$.
3.2 KAM tori analysis
We note that the Kolmogorov–Arnold–Moser (KAM) theory [46] describes how invariant tori exists in nearly integrable Hamiltonian systems. In typical cases, these invariant tori break up through destruction of homoclinic orbits (so-called homoclinic tangles) leading to chaos, which is known as torus breakdown [31, 45]. However, before we can qualify the nature of the chaos in our system as torus breakdown (seems possible from the numerical solutions from the previous section), we need some more clarifications in this regard.
3.2.1 Preliminary analysis
Let us write down our system represented by Eq.(32) in a generalized form
$\ddot{x}A(x)+\dot{x}^{2}B(x)+C_{0}(x)+\epsilon f(\tau)=0,$ (37)
where $x=v_{i}$ and the coefficient $C(x,\tau)$ is broken up into an autonomous and a time-dependent part
$C_{0}(x)$ $=$ $\frac{\mu_{0}}{(v_{d}-x)}-e^{\phi(x)},$ (38)
$\epsilon f(\tau)$ $=$ $\rho_{{\rm ext}}(\tau).$ (39)
Here, we have treated the external charged debris as a perturbation with a smallness parameter $\epsilon$. The equivalent unperturbed planar system is given by
$\begin{array}[]{rclcl}\dot{x}&=&y,\\ \dot{y}&=&F_{0}(x,y)&=&{\displaystyle\frac{-y^{2}B(x)-C_{0}(x)}{A(x)}}.\end{array}$ (40)
The equilibrium of the above system implies that $y=C_{0}(x)=0$. So, the linearized Jacobian of the vector field $(y,F_{0})$ at a point $(x_{s},y_{s})\equiv(x_{s},0)$ in equilibrium is given by
$J=\left(\begin{array}[]{cc}0&1\\ J_{21}&0\end{array}\right),\quad J_{21}=-\frac{C_{0}^{\prime}(x_{s})}{A(x_{s})},$ (41)
Figure 10: The frequency map signifying the transition to chaos through resonance.
 with the eigenvalues $\varrho$ of the Jacobian as $\varrho=\pm\sqrt{J_{21}}.$ The equilibrium point $x_{s}$ can be determined from the condition $C_{0}(x)=0$, and it comes out to be $x_{s}\simeq 1.05935$, which shows that $J_{21}<0$ and the existence of a saddle point and a homoclinic orbit can be ruled out for the unperturbed system [47]. Apparently Eq.(32) is a highly nonlinear system and as it contains a $\dot{v}_{i}^{2}$ term, it is not likely to be reducible to a Hamiltonian system as well, even in the extended phase space $(v_{i},\dot{v}_{i},\tau)$. This is also obvious from the fact that the flow is ‘non volume-preserving’ i.e. $\nabla\cdot\dot{\bm{x}}\neq 0$, where $\bm{x}=(x,y)$, which can be readily found out by observing that
$\nabla\cdot\dot{\bm{x}}\equiv\frac{\partial\dot{x}}{\partial x}+\frac{\partial\dot{y}}{\partial y}=-2y\frac{B(x)}{A(x)}\neq 0.$ (42)
All these facts indicate that a typical breakdown of KAM tori is not possible in our system. However, strong signature of torus breakdown suggests that torus breakdown may still occur, though not through tangles but via shear and resonance [48, 49, 50]. This insight leads us to carry out an analysis through frequency and finite-time Lyapunov exponent (FTLE) maps [51, 52] in the next subsection. The above analysis also shows that the system does not preserve phase space volume and dynamically acts like a dissipative system though physically it does not have any dissipation.
Figure 11: The FTLE maps for positive forcing (left) and negative forcing (right).
3.2.2 Frequency and FTLE maps
We now express our external charged debris term as a perturbation, as mentioned before $\rho_{{\rm ext}}(\tau)=\epsilon f(\tau)$, where $\epsilon$ is small but $\gtrless 0$, depending on the nature of charged perturbation. In order to compute the frequency map, we consider the actual time evolution of $x(\tau)$, as obtained from the numerical solution of Eq.(37) which is always real. From this, we construct the analytic (or complex) signal $z(\tau)$ by incorporating the Hilbert transform ${\cal H}[x(\tau)]$ of $x(\tau)$,
$z(\tau)=x(\tau)+i{\cal H}[x(\tau)],$ (43)
This helps us compute the instantaneous frequency $\omega(\tau)$
$\omega(\tau)=\frac{d}{d\tau}\theta(\tau),\quad\textrm{where}\,\theta(\tau)=\arg[z(\tau)].$ (44)
We however use the average instantaneous frequency $\bar{\omega}$, averaged over a certain time $T$ to cancel out the transient fluctuations
[…]
In Fig.10, we show a plot of $(\bar{\omega},\epsilon)$ with $\epsilon>0$ (positive charge perturbation) for a region of torus breakdown, shown in Fig.8. The figure shows that in the region $\epsilon\lesssim 0.2$ and $\gtrsim 0.4$, the instantaneous average frequency varies smoothly signifying quasi-periodic motion. In the region $\epsilon\simeq 0.25-0.35$, we have a sharp transition as well as some fluctuations indicating the onset of chaotic motion, which indicates a torus breakdown with possible resonance.
In order to compute the FTLE maps, we compute the Lyapunov exponent by calculating the divergence of orbits over a finite time window $T$, giving the FTLE [51, 52]
$\lambda_{T}(\bm{x}_{0})=\frac{1}{T}\log\frac{\left\|\delta(T)\right\|}{\left\|\delta(0)\right\|},$ (46)
where $\delta(0)$ is the initial separation of two nearby trajectories and $\delta(T)$ is the separation of the orbits at time $T$. We note that the maximal Lyapunov exponent can be recovered through $\lambda_{{\rm max}}=\lim_{T\to\infty}\lambda_{T}(\bm{x}_{0})$. The two FTLE maps – one for positive charged perturbation and the other for negative, are shown in Fig.11. As we can see that for positive charge perturbation (positive forcing with $\epsilon=+0.1$), there are dark basins with very low FTLEs on either sides with surrounding yellow-orange colored high-FTLE region. This sharp contrast is a definitive signature for large lobes of mixing of tori, indicating a torus breakdown. On the other hand, negative charge perturbation (negative forcing with $\epsilon=-0.1$) does produce chaotic regions but as we can see from the right panel, the map is more or less uniform. The FTLE maps further confirm our earlier speculation that a positive charged perturbation produces very pronounced fractal scattering indicating stronger chaos while a negative charge perturbation of the same amplitude produces a milder chaos.
Figure 12: The FTLE difference maps showing effect of stochastic fluctuation for positive forcing (left) and negative forcing (right).
3.2.3 Perturbation-dependent chaos
It is now amply clear from the numerical as well as theoretical analyses that the chaotic behavior of the plasma system with external charged debris is highly dependent on the nature of the charged debris (positive and negative). Physically, we can perhaps understand it by rewriting Eq.(32) as
$A(v_{i})\ddot{v}_{i}=-\dot{v}_{i}^{2}B(v_{i})-C(v_{i},\tau),$ (47)
which represents the total force of the system with $A(v_{i})$ acting as a mass. As $C(v_{i},\tau)=C_{0}(v_{i})+\rho_{{\rm ext}}(\tau)$, total forcing is larger when $\rho_{{\rm ext}}(\tau)>0$, while the forcing is reduced when $\rho_{{\rm ext}}(\tau)<0$ with a possibility of even cancelling the force rendering the system to a force-free state. So, positive charge perturbation pushes the system away from an equilibrium which results in a stronger chaotic state.
3.2.4 Effect of stochastic fluctuations
In practice, we would generally expect a stochastic component to the fluctuating charge density of the external debris, which can be modeled as a noise to the $\rho_{{\rm ext}}(\tau)$
$\rho_{{\rm ext}}(\tau)=\rho_{{\rm det}}(\tau)+\epsilon_{{\rm noise}}\xi(\tau),$ (48)
where $\epsilon_{{\rm noise}}$ denotes the strength of the noise (or stochastic part) while the first term denotes the deterministic part. In Fig.12, we show the FTLE difference maps for two cases $\epsilon=\pm 0.1$ where the difference between the usual FTLE and one with stochastic component is shown. We can see from the figure that in both cases $(\epsilon=\pm 0.1)$, the noise tends to destabilize weakly regular (non-chaotic or marginally chaotic) areas, whereas highly chaotic zones are least affected. So, stochastic forcing primarily affects the marginal or near-integrable areas and overall chaotic property of the system remains broadly same, a result which is confirmed from our PIC simulation.
Figure 13: The ion energy spectrum $|E_{\omega}|\sim\left(v_{i}^{2}\right)$ for both negatively and positively charged debris with and without fluctuation. The upper blue colored panels (a) and (b) are for negatively and positively charged debris, respectively without any periodic charge fluctuation of the debris. The lower red colored panels (c) and (d) are for negatively and positively charged debris with periodic charge fluctuation of the debris.
4 Spectral signature of the phase space dynamics
In this section, we provide an analysis on the basis of spectral signatures of the chaotic phase space dynamics from the PIC simulation data. The analysis broadly confirms our earlier observations given in Sections 2 and 3. In Fig.13, we have shown the ion energy spectrum $|E_{\omega}|\sim\left(v_{i}^{2}\right)$ for both negatively and positively charged debris with and without debris charge fluctuation. The first observation is that the energy spectra becomes flattened at higher frequencies which points to the energy redistribution among the waves and particles due to linear and NLLD [29, 53, 54]. We note that the system has no physical dissipation which would have steepened the spectra at the right ends indicating loss of energy. However, in absence of any physical dissipation, the Landau damping merely redistributes the energy keeping the total energy content of the system to remain constant, which causes this flattening.
The other important observation is the broadening of the spectrum in presence of periodic charge fluctuation of the debris. This broadening indicates that the coherent cascading of energy is no longer possible when there is a periodic charge fluctuation of the debris. This also points toward the limitation imposed by chaotic oscillations on coherent transfer of energy from wave to particle and vice versa. The higher broadening of the spectrum for fluctuation of positively charged debris [panel (d)] indicates more loss of coherence due to a stronger chaos [55, 56]. These energy spectra also suggest that the nonlinear structures produced due to negatively charged debris are dynamically more robust and are less prone to any charge fluctuation of the debris.
The spike at $\omega\sim 43\omega_{pi}$ in the spectra is due to the electron plasma frequency which basically shows that the primary driven mechanism of the ion oscillation is due to the electron response.
5 Conclusions
In this work, we have looked into the dynamics governing the interplay between external charged debris and a flowing $e$-$i$ plasma with a particular emphasis on the role of the periodic fluctuation of debris charge. Our analysis is focused on how these fluctuations, introduced as an external forcing in our model, affect the plasma response. We note that periodic charge fluctuation on external debris can occur naturally due to a feedback response to the IAW induced by the presence of the external debris. Our analysis shows emergence of complex nonlinear phenomena including the emergence of chaos and the manifestation of NLLD.
We took a dual simulation approach to unravel these dynamics using a kinetic particle-in-cell (PIC) simulation and flux-corrected transport-based (FCT) fluid simulation. Our findings reveal that the charge polarity of the debris plays a crucial role in determining the system’s behaviour. The charge fluctuation of the debris is implemented through a periodic external forcing model in both the simulations with the charge polarity fixed as either positive or negative throughout the simulations. Some key findings of our simulations are:
  • 1.
    The debris charge fluctuation acts as the trigger for chaotic oscillations, regardless of the charge polarity and the relative velocity between the plasma and debris.
  • 2.
    For a positively charged debris, we have observed a much stronger chaotic response confined at the debris site, with the maximal Lyapunov exponent increasing with the debris velocity. In contrast, negatively charged debris creates a more uniform weak chaos pattern across the domain.
  • 3.
    When the frequency of debris velocity approaches IA speed $v_{d}\sim c_{i}$, both PIC and FCT simulations detect damping of the nonlinear wave, which can be due to both nonlinear wave-particle resonance (NLLD) as well as nonlinear wave-wave interactions. However, while FCT simulation can only show wave-wave interactions, PIC simulation can show both the effects. The existence of NLLD is validated from the flattening of the ion velocity distribution.
  • 4.
    The added stochastic component to the regular periodic fluctuation does not affect the chaotic ion-acoustic oscillations significantly.
Theoretically, we have shown that the emergence of chaos due to the periodic charge fluctuation of the debris is through torus breakdown, though the breakdown happens via shear and resonance [48, 49, 50], not through the usual process of destruction of homoclinic orbits. The primary results of the theoretical analysis are:
  • 1.
    The periodic fluctuation of debris charge in a flowing plasma leads to chaotic ion-acoustic oscillations and the chaos occurs through a phase of torus breakdown. This is substantiated by a KAM tori analysis.
  • 2.
    The reason why a periodically fluctuating negative debris charge induces a much weaker chaos than positive debris charge is due to the fluctuation forcing the system away from a dynamically stable configuration when the debris charge is positive.
  • 3.
    Through the FTLE maps, the theoretical analysis has also been able to show why the stochastic component affect the chaos only negligibly.
At the end, we have carried out an energy spectrum analysis of the chaotic phase space, which further confirms the dynamic stability of the nonlinear structures induced by negatively charged debris. Our work adds a new aspect to the science of the detection of LEO space debris. While existing methods leverage on plasma oscillations induced by the charged debris, the different chaotic signatures for positive and negatively charged debris as we have observed here add another layer to the detection scenario which can make the management of space debris even more robust.